Consider two large, plane mirrors placed upright and making an angle of 62.4 degrees with each other. If a horizontal ray of light is incident on one mirror at an angle of incidence of 30.6 degrees, what is the angle of reflection for the ray after it is reflected from both mirrors? Calculate the answer in degrees and rounded to three significant figures.

Do not round until the end
Answer #1

卜2

i1 = 30.6 degree

since

angle of incidence = reflection,

r1 = 30.6 degree

x = 90 - r1

= 90 - 30.6

= 59.4 degree

now in triangle:

x + y + 62.4 = 180

59.4 + y + 62.4 = 180

y = 58.2 degree

i2 = 90-y1

= 90 - 58.2

= 31.8

since

angle of incidence = reflection,

r2 = 31.8 degree

Answer: 31.8 degree